# LA-MARRVEL: A Knowledge-Grounded and Language-Aware LLM Reranker for AI-MARRVEL in Rare Disease Diagnosis

### v1

@@ Introduction
Rare diseases affect millions of people1, yet each single condition is uncommon and often hard to diagnose. Many patients wait 4.7 years for an answer2. Modern DNA and RNA tests can identify genes that may explain a person’s symptoms, but they typically return a long list of candidates 3, 4. Clinicians then have to verify across papers, databases, and case reports to decide which genes truly match the patient. This step is labor-intensive and time-consuming, and it is easy to overlook critical details embedded within plain language5.
@@
Large language models (LLMs) can read and explain complex medical text. For example, the recent models show higher than 95% of accuracy for a USMLE (United States Medical Licensing Examination)-style benchmark 6, 7. However, we previously showed that LLMs may exhibit bias toward well-studied genes and sensitivity to the ordering of candidate genes, which can compromise the reliability of gene-ranking results 8. These limitations suggest that, for clinical decision-support tasks, LLMs requires grounding and stability: linking outputs to citable knowledge (grounding) and producing consistent responses across repeated runs (stability). Without these properties, even strong language skills are not enough for safe clinical use 9, 10. A major reason for this challenge is that medical knowledge is largely written as free text; for example, about 80% of all clinical data are found in this unstructured form 11. Even when symptoms and inheritance are encoded using ontologies, converting them into simple tabular formats for traditional machine learning still demands extensive normalization and feature engineering12 or loses important context. The computational gene-prioritization tools for rare-disease diagnosis are typically used as a high-recall first stage to filter/rank an initial list of candidate genes, identifying many plausible targets but still leaving a few false positives
citepJacobsen2022. Turning that a long gene list into a short, trustworthy set of likely genes requires careful reading of unstructured text and cross-checking sources —the kind of labor-intensive task that diverts valuable time and attention from busy clinical practices13, 14.
@@
To bridge these gaps, we introduce LA-MARRVEL, a language-aware reranking system that sits on top of a knowledge-driven high-recall pipeline. Rather than supplanting established bioinformatics tools, our goal is to further prioritize candidate genes, yielding a more concise, accurate, and defensible list. LA-MARRVEL brings together three pieces. First, we provide expert-engineered context: curated phenotype and disease information that gives the model the right facts at the right time. Second, we use a ranked voting method (Tideman’s method)15 to combine multiple responses into a single, consensus gene order. The ranked voting method decreases randomness and favors choices supported by repeated evidence. Lastly, we integrate this layer with the AI-MARRVEL pipeline 16, which already produces strong first-stage ranks and gene annotations. (Figure˜1)
@@
We evaluate LA-MARRVEL on three real-world patient cohorts —Baylor Genetics (BG), Deciphering Developmental Disorders (DDD), and the Undiagnosed Diseases Network (UDN). These datasets come from independent cohorts which are representative of cases commonly encountered in clinical practice. We compare our approach to widely used diagnostic tools (e.g., Exomiser17 and LIRICAL18) and to general-purpose LLMs. Across the three cohorts, we demonstrated LA-MARRVEL improves top-rank accuracy without compromising recall, resulting in more reliable prioritization of the causal gene and consequently reducing the downstream review burden.
@@ Results
  • Strong rescue at poorer ranks. For BG the bars for original ranks 3-9 are entirely green (100% improved; small $n$), and for DDD and UDN most ranks $\geq 3$ are also 100% improved. That indicates LA-MARRVEL is very effective at promoting correct candidates that the baseline ranked lower.
@@
LLMs Require a First-Stage Ranker for Effective Prioritization.
Directly prompting state-of-the-art LLMs to name causal genes performs poorly compared with rank-then-rerank pipelines. In our evaluation (Figure˜4), prompt-only LLMs (Claude and Claude-Thinking, both without AIM) recover only ~12-15% of cases at Recall@1, far below Exomiser (~50%), LIRICAL (~31%), and LA-MARRVEL (~78%). The gap persists at wider review windows: by Top-5/Top-10, LLM-only recall remains ~37-55% versus ~75-85% for classical tools and ~95-98% for LA-MARRVEL; even at Top-100 they plateau around ~70-76% compared with ~96-99% for the other methods.
These results indicate that LLMs are not reliable stand-alone retrieval engines for rare-disease gene discovery, when they are given the exhaustive list of all variant-bearing genes for each patient (benign and pathogenic alike). Faced with a long-tail of gene–phenotype links and a noisy, unfiltered candidate set, they miss true genes at clinically useful top-k cutoffs; free-form generation adds off-list distractors and hallucinated rationales; and reasoning alone ("Thinking") yields only modest gains. By contrast, pairing an LLM with a high-recall first-stage ranker (AI-MARRVEL) focuses the model on a comprehensive, evidence-grounded candidate set, enabling phenotype-aware judgment where it excels. Consequently, LA-MARRVEL’s rank-then-rerank design is essential for achieving high precision at the top ranks while maintaining near-ceiling recall.
@@
We observed that removing phenotype knowledge (Human Phenotype Ontology, HPO)19 from prompts causes the largest drop in retrieval accuracy, especially at the clinically crucial top ranks. At Top-1, the Recall@1 loss for Without HPO Information is 20.22%p, far exceeding Without Disease Information (5.90%p), Without Disease Phenotypes But with Name (4.78%p), and Without HPO Text Description But with ID (3.37%p). The pattern holds across $K$: at Recall@3, No HPO loses 10.39%p versus 1.97%p, 1.12%p, and 0.56%p for the other ablations, respectively. In short, removing phenotype knowledge produces the single largest degradation in performance. (Figure˜5)
@@
  • HPO IDs carry most of the value: When HPO IDs are retained but their human-readable text is removed (No HPO Text), losses are small (e.g., 3.37%p at K=1), indicating the HPO ID itself is highly informative.
@@
Increasing the number of candidate genes ($G$) shifts retrieval behavior in two opposing ways. With fewer candidates, the model is more decisive and more often places the correct gene at rank 1 boosting Recall@1. As $G$ grows, the pool broadens and the system captures more true positives within the top 10 raising Recall@10. In our experiments, ($G$ = 20, 30, 50, 100) consistently improved Recall@10 over the first-stage ranker (AI-MARRVEL) simply because more plausible genes were evaluated. (Figure˜6)
@@
Across all gene-set sizes ($G$ = 10, 20, 30, 50, 100) and for every K, Ranked Voting (10x) yields higher Recall@K than a Single Run (1x). The advantage is visible from small K (Top-2/Top-3) through larger K, and the curves do not cross —Ranked Voting — i.e., always perform better than Single Run. (Figure˜7)
@@
To facilitate the review of gene rankings, LA-MARRVEL generates an explainer trace — a step-by-step view that adds evidence to each candidate gene. Each panel shows ACMG evidence codes — the American College of Medical Genetics and Genomics criteria for classifying pathogenic variants20 — zygosity checks, and how well the phenotype matches HPO terms. This trace follows the manual curation process and records the reasons for any changes in rank.
@@
In this case, AI-MARRVEL put CLDN16 at rank 1 and SPG7 at rank 2, even though the true diagnosis is linked to SPG7 (Figure˜8). LA-MARRVEL’s reasoning trace walked through the same evidence a human would check and corrected the order.
@@
In this case, AI-MARRVEL initially ranked UNC80 first and STIM1 ninth, even though the ground truth is linked to STIM1 (Figure˜9). LA-MARRVEL’s explainable trace re-evaluated the molecular, inheritance, and phenotype evidence and corrected the ranking. “‘
@@ Discussion
  • DDD (Deciphering Developmental Disorders). 214 cases; on average 811.5 candidate genes, 7.38 HPO terms, and 1.000 causal genes per case, spanning 116 unique causal genes. Curated by AMELIE; accessibility is Restricted.
@@
Hyperparameters Unless otherwise specified, all decoding and sampling hyperparameters (e.g., temperature, top_k, top_p, and related settings) were kept at their default values provided by the Bedrock-hosted Claude-4-Sonnet endpoint. No additional tuning of these hyperparameters was performed across experiments.
@@
LLMs are evolving, but at the current stage they require concise, informative context provided via a prompt. On average, a proband has variants in 800-1,500 genes, including benign ones; including all genes in a prompt is not feasible. To reduce the number of candidate genes to a feasible size, we chose the reputable tool AI-MARRVEL, known for its state-of-the-art methods in rare disease diagnosis on the BG, DDD, and UDN cohorts 16.
@@
AI-MARRVEL not only ranks candidate genes, it also generates variant- and gene-level annotations using Ensembl’s Variant Effect Predictor (VEP) and the MARRVEL database. LA-MARRVEL then uses these to create concise gene summaries (See the below prompt template, Table˜2, and Table˜3), which can help preliminarily assess several ACMG/AMP criteria 20, including: variant consequence (PVS1, when predicted loss-of-function is a known disease mechanism); computational predictions such as CADD (PP3/BP4); gnomAD allele frequency (PM2/BA1, absent or rare / too common); gene-disease information (e.g., OMIM) combined with a highly specific patient phenotype (PP4); ClinVar (PP5/BP6, reputable database).
@@
Ensembling LLMs at inference time —via self-consistency (sample diverse chains of thought and pick the modal answer), structured search like Tree-of-Thoughts, or Best-of-$N$ selection with a verifier —consistently boosts accuracy on math, code, and reasoning benchmarks 21, 22, 23. In addition, sampling-efficient variants such as ConSol use Sequential Probability Ratio Testing (SPRT) to adaptively stop once responses converge, cutting the number of samples required relative to vanilla self-consistency while preserving accuracy 24.
Here, each sample yields a partial ranking of candidate genes (unranked items are treated as missing). We aggregate these ballots using Ranked Pairs (Tideman) 15: (1) build pairwise win counts $\text{Wins}[a,b]$ by treating any ranked item as preferred over any candidate genes; (2) for each unordered pair, form a directed victory with margin $|\text{Wins}[a,b]-\text{Wins}[b,a]|$; (3) sort directed pairs by decreasing margin, then by the winner’s votes, with deterministic lexicographic tie-breaks; (4) greedily lock edges that do not create cycles; (5) output a topological order, breaking ties by Borda score (computed only from ranked positions) then by name (see Algorithm˜1 and 2). Ranked Pairs is Condorcet-compliant and clone-independent, leveraging full list structure while remaining robust to near-duplicate options.

### v2

@@ Introduction
Rare diseases affect millions of peopleTheLancetGlobalHealth2024, yet each single condition is uncommon and often hard to diagnose. Many patients wait 4.7 years for an answerFaye2024. Modern DNA and RNA tests can identify genes that may explain a person’s symptoms, but they typically return a long list of candidates IOM1994GeneticRisks, Murdock2020. Clinicians then have to verify across papers, databases, and case reports to decide which genes truly match the patient. This step is labor-intensive and time-consuming, and it is easy to overlook critical details embedded within plain languageBirgmeier2020.
@@
Large language models (LLMs) can read and explain complex medical texts. For example, the recent models show higher than 95% accuracy for a USMLE (United States Medical Licensing Examination)-style benchmark Jin2021, ValsMedQA2025. However, we previously showed that LLMs may exhibit bias toward well-studied genes and sensitivity to the ordering of candidate genes, which can compromise the reliability of gene-ranking results Neeley2024. These limitations suggest that, for clinical decision-support tasks, LLMs require grounding and stability: linking outputs to citable knowledge (grounding) and producing consistent responses across repeated runs (stability). Without these properties, even strong language skills are not enough for safe clinical use Wu2025, Asgari2025. A major reason for this challenge is that medical knowledge is largely written as free text; a large proportion of clinical data is found in this unstructured form Murdoch2013. Even when symptoms and inheritance are encoded using ontologies, converting them into simple tabular formats for traditional machine learning still demands extensive normalization and feature engineeringGarla2012 or loses important context. The computational gene-prioritization tools for rare-disease diagnosis are typically used as a high-recall first stage to filter/rank an initial list of candidate genes, identifying many plausible targets but still leaving a few false positivesJacobsen2022. Turning a long gene list into a short, trustworthy set of likely genes requires careful reading of unstructured text and cross-checking sources —the kind of labor-intensive task that diverts valuable time and attention from busy clinical practicesBrownstein2014, Dewey2014.
@@
To bridge these gaps, we introduce LA-MARRVEL, a language-aware reranking system that sits on top of a knowledge-driven high-recall pipeline. Rather than supplanting established bioinformatics tools, our goal is to further prioritize candidate genes, yielding a more concise, accurate, and defensible list. LA-MARRVEL brings together three pieces. First, we provide expert-engineered context: curated phenotype and disease information that gives the model the right facts at the right time. Second, we use a ranked voting method (Tideman’s method)Tideman1987 to combine multiple responses into a single, consensus gene order. The ranked voting method decreases randomness and favors choices supported by repeated evidence. Lastly, we integrate this layer with the AI-MARRVEL pipeline Mao2024, which already produces strong first-stage ranks and gene annotations. (Figure˜1)
@@
We evaluate LA-MARRVEL on three real-world patient cohorts —Baylor Genetics (BG), Deciphering Developmental Disorders (DDD), and the Undiagnosed Diseases Network (UDN). These datasets come from independent cohorts which are representative of cases commonly encountered in clinical practice. We compare our approach to widely used diagnostic tools (e.g., ExomiserJacobsen2022 and LIRICALRobinson2020) and to general-purpose LLMs. Across the three cohorts, we demonstrated LA-MARRVEL improves top-rank accuracy without compromising recall, resulting in more reliable prioritization of the causal gene and consequently reducing the downstream review burden.
@@ Results
  • Strong rescue at poorer ranks. For BG the bars for original ranks 3–9 are entirely green (100% improved; small $n$), and for DDD and UDN most ranks $\geq 3$ are also 100% improved. That indicates LA-MARRVEL is very effective at promoting correct candidates that the baseline ranked lower.
@@
We observed that removing phenotype knowledge (Human Phenotype Ontology, HPO)Khler2020 from prompts causes the largest drop in retrieval accuracy, especially at the clinically crucial top ranks. At Top-1, the Recall@1 loss for Without HPO Information is 20.22 percentage point, far exceeding Without Disease Information (5.90 points), Without Disease Phenotypes But with Name (4.78 points), and Without HPO Text Description But with ID (3.37 points). The pattern holds across $K$: at Recall@3, Without HPO Information loses 10.39 points versus 1.97 points, 1.12 points, and 0.56 points for the other ablations, respectively. In short, removing phenotype knowledge produces the single largest degradation in performance. (Figure˜4)
@@
  • HPO IDs carry most of the value: When HPO IDs are retained but their human-readable text is removed (Without HPO Text Description But with ID), losses are small (e.g., 3.37%p at K=1), indicating the HPO ID itself is highly informative.
@@
Increasing the number of candidate genes ($G$) shifts retrieval behavior in two opposing ways. With fewer candidates, the model is more decisive and more often places the correct gene at rank 1 boosting Recall@1. As $G$ grows, the pool broadens and the system captures more true positives within the top 10 raising Recall@10. In our experiments, ($G$ = 20, 30, 50, 100) consistently improved Recall@10 over the first-stage ranker (AI-MARRVEL) simply because more plausible genes were evaluated. (Figure˜5)
@@
Across all gene-set sizes ($G$ = 10, 20, 30, 50, 100) and for every K, Ranked Voting (10x) yields higher Recall@K than a Single Run (1x). The advantage is visible from small K (Top-2/Top-3) through larger K, and the curves do not cross —Ranked Voting — i.e., always perform better than Single Run. (Figure˜6)
@@
To facilitate the review of gene rankings, LA-MARRVEL generates an explainer trace — a step-by-step view that adds evidence to each candidate gene. Each panel shows ACMG evidence codes — the American College of Medical Genetics and Genomics criteria for classifying pathogenic variantsRichards2015 — zygosity checks, and how well the phenotype matches HPO terms. This trace follows the manual curation process and records the reasons for any changes in rank.
@@
In this case, AI-MARRVEL put CLDN16HGNC_CLDN16 at rank 1 and SPG7HGNC_SPG7 at rank 2, even though the true diagnosis is linked to SPG7 (Figure˜7). LA-MARRVEL’s reasoning trace walked through the same evidence a human would check and corrected the order.
@@
In this case, AI-MARRVEL initially ranked UNC80HGNC_UNC80 first and STIM1HGNC_STIM1 ninth, even though the ground truth is linked to STIM1 (Figure˜8). LA-MARRVEL’s explainable trace re-evaluated the molecular, inheritance, and phenotype evidence and corrected the ranking. “‘
@@ Discussion
  • DDD (Deciphering Developmental Disorders). 214 cases; on average 811.5 candidate genes, 7.38 HPO terms, and 1.000 causal genes per case, spanning 116 unique causal genes. Curated by AMELIEBirgmeier2020; accessibility is Restricted.
@@
Hyperparameters Unless otherwise specified, all decoding and sampling hyperparameters (e.g., temperature, top_k, top_p, and related settings) were kept at their default values provided by the Bedrock-hosted Claude-4-Sonnet endpointAWS_CLAUDEPARAMS. No additional tuning of these hyperparameters was performed across experiments.
@@
LLMs are evolving, but at the current stage they require concise, informative context provided via a prompt. On average, a proband has variants in 800–1,500 genes, including benign ones; including all genes in a prompt is not feasible. To reduce the number of candidate genes to a feasible size, we chose the reputable tool AI-MARRVEL, known for its state-of-the-art methods in rare disease diagnosis on the BG, DDD, and UDN cohorts Mao2024.
@@
AI-MARRVEL not only ranks candidate genes, it also generates variant- and gene-level annotations using Ensembl’s Variant Effect Predictor (VEP) and the MARRVEL database. LA-MARRVEL then uses these to create concise gene summaries (See the below prompt template, Table˜2, and Table˜3), which can help preliminarily assess several ACMG/AMP criteria Richards2015, including: variant consequence (PVS1, when predicted loss-of-function is a known disease mechanism); computational predictions such as CADD (PP3/BP4); gnomAD allele frequency (PM2/BA1, absent or rare / too common); gene-disease information (e.g., OMIM) combined with a highly specific patient phenotype (PP4); ClinVar (PP5/BP6, reputable database).
@@
Ensembling LLMs at inference time —via self-consistency (sample diverse chains of thought and pick the modal answer), structured search like Tree-of-Thoughts, or Best-of-$N$ selection with a verifier —consistently boosts accuracy on math, code, and reasoning benchmarks WangSelfCons2023, YaoTreeThoughts2023, WangMathShepherd2024. In addition, sampling-efficient variants such as ConSol use Sequential Probability Ratio Testing (SPRT) to adaptively stop once responses converge, cutting the number of samples required relative to vanilla self-consistency while preserving accuracy LeeConSol2025.
Here, each sample yields a partial ranking of candidate genes (unranked items are treated as missing). We aggregate these ballots using Ranked Pairs (Tideman) Tideman1987: (1) build pairwise win counts $\text{Wins}[a,b]$ by treating any ranked item as preferred over any candidate genes; (2) for each unordered pair, form a directed victory with margin $|\text{Wins}[a,b]-\text{Wins}[b,a]|$; (3) sort directed pairs by decreasing margin, then by the winner’s votes, with deterministic lexicographic tie-breaks; (4) greedily lock edges that do not create cycles; (5) output a topological order, breaking ties by Borda score (computed only from ranked positions) then by name (see Algorithm˜1 and 2). Ranked Pairs is Condorcet-compliant and clone-independent, leveraging full list structure while remaining robust to near-duplicate options.
@@ Acknowledgments
Algorithm 1 TidemanRankedPairs(ballots)
1:ballots: list of partial rankings over candidates (1 = best). Missing rank = unranked.
2:order: candidates from best to worst
3:if ballots is empty then return empty list
[…]
6:
[…]
8:initialize $\text{Borda}[c]\leftarrow 0$ for all $c\in C$
9:for each ballot $s$ do
10:  let $r$ be $s$ reindexed to $C$ (candidates absent in $s$ are unranked in $r$)
11:  for each $c\in C$ do
12:    if $r[c]$ is ranked then
13:     $\text{Borda}[c]\leftarrow\text{Borda}[c]+(m-r[c]+1)$
14:    end if
15:  end for
16:end for
17:
[…]
19:initialize $\text{Wins}[a][b]\leftarrow 0$ for all $a,b\in C$
20:for each ballot $s$ do
21:  let $r$ be $s$ reindexed to $C$
22:  for each ordered pair $(a,b)$ with $a\neq b$ do
23:    if $r[a]$ ranked and $r[b]$ unranked then
24:     $\text{Wins}[a][b]\leftarrow\text{Wins}[a][b]+1$
25:    else if $r[a]$ ranked and $r[b]$ ranked and $r[a]<r[b]$ then
26:     $\text{Wins}[a][b]\leftarrow\text{Wins}[a][b]+1$
[…]
28:  end for
29:end for
30:
[…]
33:for each unordered pair $\{a,b\}$ with $a\neq b$ and $a<b$ (lexicographic) do
34:  $w_{ab}\leftarrow\text{Wins}[a][b]$; $w_{ba}\leftarrow\text{Wins}[b][a]$
35:  if $w_{ab}>w_{ba}$ then
36:    append $(a,b,\;w_{ab}-w_{ba},\;w_{ab},\;w_{ba})$ to $P$
37:  else if $w_{ba}>w_{ab}$ then
38:    append $(b,a,\;w_{ba}-w_{ab},\;w_{ba},\;w_{ab})$ to $P$
39:  end if
40:end for
41:sort $P$ in descending order by:
42:    (i) victory margin, (ii) winner’s votes, then ascending by (iii) winner name, (iv) loser name
43:
[…]
45:initialize directed adjacency sets $\text{Adj}[c]\leftarrow\emptyset$ for all $c\in C$
46:function CreatesCycle($u,v$)
47:  return (there exists a path $v\leadsto u$ in graph Adj) $\triangleright$ e.g., DFS from $v$ following Adj edges
48:end function
49:for each $(u,v,\cdot)$ in $P$ in order do
50:  if not CreatesCycle($u,v$) then
51:    add edge $u\to v$ to Adj
52:  end if
53:end for
54:
[…]
56:compute $\text{InDeg}[c]$ from Adj
[…]
59:while Avail not empty do
60:  remove first $u$ from Avail; append $u$ to Order
61:  for each $v$ in $\text{Adj}[u]$ (any order) do
62:    $\text{InDeg}[v]\leftarrow\text{InDeg}[v]-1$
63:    if $\text{InDeg}[v]=0$ then
64:     insert $v$ into Avail
65:    end if
66:  end for
67:  resort Avail by descending Borda, then by name
68:end while
69:return Order
Algorithm 2 ReorderListByConsensus(PrioritizedLists, InitialList)
1:PrioritizedLists: outputs of LLM; InitialList: Top-G genes ordered by the score of AI-MARRVEL
2:ReorderedList: reranked gene list in consensus order, then any stragglers
[…]
5:for $\text{PrioritizedList}\in\textit{PrioritizedLists}$ do
6:  $\text{Head}\leftarrow$ items of InitialList which is in PrioritizedList, ordered by PrioritizedList
7:  $\text{Tail}\leftarrow$ items of InitialList which is not in PrioritizedList (ordered by InitialList)
8:  $\text{ballot}\leftarrow$ concatenate Head then Tail
9:  append ballot to ballots
10:end for
11:
[…]
17:
18:return ReorderedList
Table 1: Stats of BG-DDD-UDN Cohort Data
BG DDD UDN
Case 63 214 90
Avg. Gene 1469.2 811.5 1072.2
Avg. Hpo 10.59 7.38 43.98
Avg. Causal Genes 1.175 1.000 1.044
Unique Causal Genes 74 116 93
Accessibility Internal Restricted Restricted
Source Baylor Genetics DDD Undiagnosed Diseases Network
Curation In-House AMELIE In-House
Table 2: HPO summary for nystagmus
ID Name Synonyms Definition
HP:0000639 Nystagmus [’Involuntary, rapid, rhythmic eye movements’] Rhythmic, involuntary oscillations of one or both eyes related to abnormality in fixation, conjugate gaze, or vestibular mechanisms.
Table 3: Gene summary for SPG7
Gene ClinVar status CADD Score AF Zygosity Trans-Heterozygous Consequence Impact Phenotype
SPG7 Conflicting interpretations of pathogenicity 27.8 0.002923 Heterozygous No missense variant MODERATE Spastic paraplegia 7, autosomal recessive. Inheritance: AD; AR. Clinical features: Nystagmus; Lower limb muscle weakness; Urinary bladder sphincter dysfunction; Waddling gait; Lower limb hypertonia; Muscle weakness; Degeneration of the lateral corticospinal tracts; Hyperreflexia; Vertical supranuclear gaze palsy; Dysphagia; Scoliosis; Babinski sign; Gait disturbance; Memory impairment; Pes cavus; Spastic ataxia; Urinary urgency; Lower limb hyperreflexia; Upper limb muscle weakness; Dysarthria; Dysdiadochokinesis; Upper limb hypertonia; Postural instability; Limb ataxia; Upper limb hyperreflexia; Supranuclear gaze palsy; Abnormality of somatosensory evoked potentials; Optic atrophy; Spastic paraplegia; Upper limb spasticity.
Figure 1: Schematic illustration of LA-MARRVEL
AI-MARRVEL first generates high-recall, variant-bearing candidate genes with annotations. LA-MARRVEL then composes knowledge-grounded prompts using HPO terms, disease and gene summaries, and variant-level evidence, queries an LLM multiple times, and aggregates the resulting partial rankings using Tideman’s ranked-pairs voting. The final output is a reranked gene list with an explainable trace that integrates phenotype match, inheritance, and ACMG-style variant assessment.
(A)   Comparison with Naive LLMs by Recall@K
(B)   Cross-Dataset Evaluation with Bioinformatics Pipelines
Figure 2: Comparative Performance Analysis of LA-MARRVEL.
(A) Barplot of Recall@K (K = 1, 5, 10, 100) comparing two prompt-only LLM settings (Claude and Claude-Thinking, both without AI-MARRVEL context) against classical phenotype-driven tools (LIRICAL, Exomiser) and LA-MARRVEL. Prompt-only LLMs recover only ~12–15% of causal genes at Recall@1 and remain below ~55% even at Recall@10, whereas classical tools reach ~50–75% and LA-MARRVEL attains ~78% at Recall@1 and ~95—98% by Recall@10, demonstrating that an LLM alone is insufficient and that coupling it to a high-recall first-stage ranker is essential for clinically useful performance.
(B) Recall@K (K=1-10) is shown for LA-MARRVEL, AI-MARRVEL, Exomiser, and LIRICAL across the BG, DDD, and UDN cohorts. LA-MARRVEL consistently achieves the highest recall at clinically salient low K (Top-1/Top-3) while maintaining near-ceiling recall by Top-10, demonstrating improved prioritization of causal genes over both the first-stage ranker and established phenotype-driven tools.
Figure 3: Ratio of Improved and Harmed Case by Original Rank
For each cohort (BG, DDD, UDN), bars show the fraction of cases in which LA-MARRVEL improved, harmed, or left unchanged the rank of the causal gene, stratified by its initial AI-MARRVEL position. LA-MARRVEL most often rescues cases where the causal gene was originally ranked lower $(\geq 3)$, while a modest fraction of harms is concentrated among cases where the baseline already ranked the causal gene at the very top.
Figure 4: Performance Gain of Context Engineering
Recall@K loss is shown when removing key phenotype and disease information from the reranker input. Removing all HPO information yields the largest loss in Recall@K, followed by removing disease phenotypes, disease names, and HPO text descriptions (while retaining their identifiers). These results highlight the critical importance of structured HPO phenotypes and disease context for optimal performance in LA-MARRVEL.
Figure 5: Performance across different Top-$G$ candidate sets
Barplots of Recall@K for LA-MARRVEL using varying numbers of first-stage candidate genes ($G=10,20,30,50,100$). Performance monotonically improves with larger candidate sets, reflecting the value of high-recall upstream ranking. Gains are most pronounced at low $K$, where small $G$ restricts recovery of causal genes.
Figure 6: Recall Curve of Performance By Differing Top G Genes and N Outputs
Recall@K curves are shown for $G=50$ and $G=100$ under two conditions: a single LA-MARRVEL reranking pass vs. a 10-run ranked-voting aggregation. Ranked voting improves stability and recall, especially at low $K$, demonstrating that multiple-run aggregation can enhance prioritization performance when many candidate genes are available.
Figure 7: The Overview of Case Study 1
Each node represents either the patient, a gene, or a phenotype group. The phenotype node contains the list of phenotypes observed in the case. Gene nodes indicate candidate genes that are reprioritized (promoted) or deprioritized (demoted) by LA-MARRVEL. Nodes directly connected to gene nodes represent known relationships (e.g., gene–disease or gene–phenotype links) supporting or weakening their candidacy. Items shown in bold indicate matches to the patient’s observed phenotypes.
Figure 8: The Overview of Case Study 2
Each node represents either the patient, a gene, or a phenotype group. The phenotype node contains the list of phenotypes observed in the case. Gene nodes indicate candidate genes that are reprioritized (promoted) or deprioritized (demoted) by LA-MARRVEL. Nodes directly connected to gene nodes represent known relationships (e.g., gene–disease or gene–phenotype links) supporting or weakening their candidacy. Items shown in bold indicate matches to the patient’s observed phenotypes.